HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

PRECISION SPECTROSCOPY OF TRANSITION FROM THE METASTABLE 2^{3}S_{1} STATE OF TO HIGH NP RYDBERG STATES

\underline{G. CLAUSEN}, S. SCHEIDEGGER, J. A. AGNER, H. SCHMUTZ, and F. MERKT, Laboratory of Physical Chemistry, ETH Zurich

The metastable He ((1s)^{1}(2s)^{1}) atom in its singlet (^{1}S_{0}) or triplet (^{3}S_{1}) states is an ideal system to perform tests of ab-initio calculations of two-electron systems that include quantum-electrodynamics and nuclear finite-size effects. The recent determination of the ionization energy of the metastable 2^{1}S_{0} state of ^{4}He [1] confirmed a discrepancy between the latest theoretical values of the Lamb shifts in low-lying electronic states of triplet helium [2] and the measured 3^{3}D ← 2 ^{3}S [3] and 3^{3}D ← 2 ^{3}P [4] transition frequencies. This discrepancy could not be resolved in the latest calculations [5, 6]. Currently, we focus on the development of a new experimental method for the determination of the ionization energy of the 2^{3}S_{1} state of ^{4}He via the measurement of transitions from the 2^{3}S_{1} state to np Rydberg states. Extrapolation of the np series yields the ionization energy with sub-MHz accuracy. In this poster, we present the progress in the development of our experimental setup, which involves (i) the preparation of a cold, supersonic expansion of helium atoms in the 2^{3}S_{1} state, (ii) the development and characterization of a laser system for driving the transitions to the np Rydberg states, and (iii) the implementation of a new sub-Doppler, background-free detection method. We present this new spectroscopic method, with which we cancel the ^{1}st-order Doppler shift and illustrate its power with a new determination of the ionization energy of 2^{3}S_{1} metastable He.

References

[1] G. Clausen et al., Phys. Rev. Lett. 127, 093001 (2021).

[2] V. Patkóš et al., Phys. Rev. A. 103, 042809 (2021).

[3] C. Dorrer et al., Phys. Rev. Lett. 78, 3658 (1997).

[4] P.-L. Luo et al., Phys. Rev. A. 94, 062507 (2016).

[5] V. A. Yerokhin et al., Eur. Phys. J. D. 76, 142 (2022).

[6] V. A. Yerokhin et al., Phys. Rev. A. 107, 012810 (2023).

p-number: p020  Submitted on Wed, 31 May 2023 15:47:06 +0200